### v2

# Floralens: A Deep Learning Model for the Portuguese Native Flora

Abstract
Machine-learning techniques, especially deep convolutional neural networks, are pivotal for image-based identification of biological species in many Citizen Science platforms. In this paper, we describe the construction of a dataset for the Portuguese native flora based on publicly available research-grade datasets, and the derivation of a high-accuracy model from it using off-the-shelf deep convolutional neural networks. We anchored the dataset in high-quality data provided by Sociedade Portuguesa de Botânica and added further sampled data from research-grade datasets available from GBIF. We find that with a careful dataset design, off-the-shelf machine-learning cloud services such as Google’s AutoML Vision produce accurate models, with results comparable to those of Pl@ntNet, a state-of-the-art citizen science platform. The best model we derived, dubbed Floralens, has been integrated into the public website of Project Biolens, where we gather models for other taxa as well. The dataset used to train the model is also publicly available on Zenodo.
Keywords: automatic identification, citizen science, deep learning, computer vision
1 Introduction
The improvements in processing speed, storage capacity, and imaging sensors for mobile devices paved the way for Citizen Science [1] applications and Web services that allow amateur enthusiasts to participate in science projects. One successful case study is nature observation, specifically the photographic recording of animals, plants, and fungi in their natural habitats. Besides storing these observations, some platforms use deep-learning models to provide automatic taxonomic identification from user-provided images [2, 3, 4, 5, 6]. The data gathered by such projects is valuable for scientists, from hardcore taxonomists to ecologists studying the impact of human activity on biodiversity [7, 8].
In project Biolens [9], we are interested in creating lightweight identification models for web and mobile apps, possibly working offline (e.g., in the field) and having the option of not immediately sharing data (e.g., for research projects). Citizen science platforms do not provide this level of flexibility. Moreover, a detailed description of their methodological approach is lacking in the literature (cf. Section 2) making it difficult to use this accumulated experience for new projects. We therefore decided to independently develop a streamlined methodology to build datasets for native Portuguese species covering different taxa and to derive accurate CNN-based models using off-the-shelf machine-learning tools. Currently, we have four models: Lepilens and Mothlens (for butterflies and moths, together covering the order Lepidoptera); Dragonlens (for dragonflies and damselflies, covering the order Odonata), and, the latest addition, Floralens (described in this paper, covering the Plantae kingdom).
Figure 1: Detail of the FloraOn web application.
This paper describes the derivation of Floralens, a high-accuracy machine-learning model for automatic taxonomic identification of the Portuguese native flora. The work is anchored on the FloraOn dataset provided by the Sociedade Portuguesa de Botânica [10], available online via a web application (Figure 1) and as a contributed dataset in the Global Biodiversity Information Facility (GBIF) [11]. This dataset contains relatively few images per species but those provided are of very high quality and identified by experienced taxonomists. We use this list of Portuguese native species as our reference to build the dataset.
The methodology used to derive all the Biolens models has two core traits: (a) the use of data from research-grade public repositories for dataset construction, and; (b) the use of Google’s AutoML Vision (GAMLV) to derive the actual models. This methodology is briefly described in a short scientific outreach article (in Portuguese) [12], and, more thoroughly, in MSc theses [13, 14] (both covering different stages of the work on Floralens), and a BSc project report [15] (covering Lepilens). Our evaluation of Floralens shows that, with a carefully designed dataset, current off-the-shelf machine-learning cloud-based services output models whose performance rivals that obtained with models provided by state-of-the-art citizen science projects.
The main contributions of this work are as follows:
  • a dataset for the Portuguese Flora based on published research-grade datasets, available from Zenodo [16];
  • a high-accuracy GAMLV-based model for the Portuguese native flora publicly available via web and mobile applications; and
  • a quantitative evaluation of the derived model and a comparison of its accuracy relative to the state-of-the-art platform Pl@ntNet.
The remainder of this paper is structured as follows. Section 2 describes the current state-of-the-art regarding automatic taxonomic identification based on deep learning. Section 3 describes the construction of the dataset used in this study. Section 4 describes the generation of the model from the dataset using GAMLV. Section 5 describes the results obtained with the model. Section 6 describes the software artifacts and datasets produced in the scope of this work. Finally, Section 7 summarizes the main findings of this study and puts forward some future research goals.
2 Related Work
Convolutional Neural Networks (CNN) are deep neural networks composed of multiple layers of trainable convolutional nodes whose aggregate outputs are eventually fed to a final fully connected layer. They are often used for image classification. In this context, a convolution is an operation that applies a matrix known as a kernel to a given input matrix (an image or part of it). The kernel slides over the input, multiplying the overlapping matrix positions at each value and then summing the values. Depending on the form of the kernel, the resulting matrix can encode features such as edges, textures, and shapes, extracted from the original image. Recently, Vision Transformers emerged as powerful tools to derive accurate models for image classification tasks [17, 18, 19].
The advent of deep learning allowed the development of the first tools for automatically identifying plant species from input images [20, 21, 22]. The success of these first efforts and their further refinement quickly reached a point in which automatic species identification rivaled identifications made by specialists [23], thus attesting to the transformative role of AI in this field [24]. Nowadays, models based on deep learning are central tools in major citizen science platforms such as iNaturalist [2], Observation.org [3] and Pl@ntNet [6] and are integrated in web and mobile applications. However, the training process and machine learning methodologies used by these citizen science platforms are not documented in detail in the literature: [2] merely makes reference to machine learning from an end-user perspective, some details can only be found online in scattered form [25]; [6] provides only a summary of the Pl@ntNet methodology to derive models, the information is also given in slightly more extended form in [26]; [3] is a short abstract concerning their mobile app ObsIdentify [27], and both the app and the site interface are enabled by an API [24, 28] whose underlying model is not explained.
Deriving a CNN-based model or other ML models often requires expert knowledge, non-trivial configuration aspects, writing specialized code, and a heavyweight computational infrastructure for storing data and model training. The strain is more acute when the amount of data involved is non-negligible, as in the case of Floralens where we use approximately 300,000 photos to derive a model. AutoML cloud services address these needs, of which GAMLV is an example. Like Google, other major cloud providers have MLaaS offerings, e.g., Amazon Rekognition [29], Apple Create ML [30], and Azure AutoML [31]. These services allow the automation of several aspects of model derivation and are backed by computational infrastructures that include special-purpose hardware such as GPUs or TPUs, necessary to derive a model in a reasonable amount of time. Models derived using these tools are referenced in the literature of several areas of knowledge, e.g., biology [32], medicine [33, 34], agriculture [35], engineering [36]. Comparative results between GAMLV and one or more competitor platforms are provided in some of these works [32, 33, 34], generally illustrating good performance by GAMLV-derived models. We should note that our choice of GAMLV was solely motivated by the availability of Google Cloud research credits, which we could not obtain from other providers.
Citizen science platforms also generate important by-products in the form of their curated datasets, often made available to the public through biodiversity data portals, notably GBIF [11]. These datasets enable the development of other ML models as is the case of Floralens that, in addition to FloraOn, uses data sampled from GBIF datasets provided by iNaturalist, Observation.org, and Pl@ntNet (cf. Section 3).
Compared to the previously mentioned citizen science platforms whose domain of application is the global flora, Floralens is more specialized covering only the Portuguese native species. It is also not supported by a citizen science platform nor, for the time being, supports directly exporting data to them. Users can only export their observations into common data formats such as CSV and ZIP for images. In this respect, Floralens is closer in philosophy and implementation to the Flora Incognita Project [5]. The projects share the goal of devising machine learning models and related applications for flora identification while falling short of integration with citizen science platforms (e.g., iNaturalist, Pl@ntNet, Observation.org) or enabling data aggregation (e.g., GBIF).
To improve the identification precision, some platforms such as Pl@ntNet developed regional models by dividing the globe into several biogeographic domains [37] following published regional floras such as the WCVP/Kew [38], and introduced metadata providing information on the anatomic part of the plant depicted in an image, e.g., flower, leaf, stem. This extra metadata critically improves the precision of the models [39, 40]. At this stage, Floralens does not use image metadata to aid in the identification. A different trend that has recently generated considerable interest is the generation of global models based on extreme datasets, consisting of millions of images representing tens of thousands of individual species (for comparison, it is estimated that the world has $\sim$300K plant species), as exemplified by iNat Challenge [41] and PlantCLEF/LifeCLEF [42, 43].
3 Dataset Construction
We now describe the creation of the Floralens dataset, subsequently used for training the deep learning model.
3.1 Data Sources
Our universe of species was the FloraOn catalog as of November 2021. It covers 2,712 species of native Portuguese flora. We started the construction of the dataset by gathering a large collection of images for each species in our list. These were extracted from curated datasets available from GBIF. We then sampled this collection so that each species would be represented by at least 50 and, at most, 200 images in the dataset. The goal was to ensure that a minimum number of images was available for each species and that a species would not be over-represented as it would introduce bias in the models derived from the dataset. These bounds were defined based on our prior experience building datasets for other biological taxa (cf. Section 6).
The FloraOn repository contains geo-referenced records of Flora species with associated images. The image data is relatively broad in scope, as it covers 78% of the entire catalog (2,127 out of 2,712 species), but has a limited volume: on average there are just 11 images per species, and, unsurprisingly, our 50-image lower bound threshold is not met for a single species. Hence, to adequately populate the Floralens dataset, we retrieved the FloraOn images but also considered image data from three publicly-available datasets stored and made available at GBIF by three citizen science platforms: iNaturalist [44], Observation.org [45], and Pl@ntNet [46].
Figure 2: The geographical region of interest for GBIF portal queries.
The GBIF datasets provide validated observation data and associated images, originally submitted by users of the respective platforms. However, the validation process differs among data sources, as discussed further in this section. For each dataset and each species in our universe, we used GBIF portal queries to obtain observation records in the Darwin Core Archive format [47]. Each such record corresponds to an observation of a specimen, typically made in the wild by citizen scientists or experts, accompanied by its taxonomic identification, its geographical location, and one or more images. The GBIF portal queries were parameterized to cover the European continent (Figure 2) as, in a preliminary analysis, we found that occurrence data from just Portugal or even the entire Iberian Peninsula would yield limited data in terms of volume and variety. Fortunately, many species of the Portuguese flora are widely distributed.
Table 1: Raw data and derived dataset after sampling (#I: image count; #S: species count; $\geq 50$ I: species with more than 50 images).
Raw data Dataset
Source #I %I #S $\geq 50$ I #I %I #S
FloraOn 22,869 0.5 2,127 0 15,191 5.1 1,397
iNaturalist 2,753,167 66.2 2,066 1,431 90,127 30.6 1,358
Observation.org 823,389 19.8 1,816 1,114 85,746 29.2 1,093
Pl@ntNet 515,950 12.4 1,495 735 102,537 34.9 1,373
Total 4,154,895 2,539 1,678 293,601 1,678
3.2 Assembly
The raw data, from all image sources, is listed in Table 1 (left), along with the characterization of the Floralens dataset (right) that results from sampling the raw data. The corresponding histograms, depicting the number of species versus the number of images, are illustrated in Figure 3. In the raw data, more than 4 million images were available for consideration, covering 2,539 species (93% of the FloraOn catalog). Only 0.5% of these images are from FloraOn, and approximately two-thirds are taken from iNaturalist. Moreover, only 1,678 species reached our lower bound threshold of 50 images (61% of the FloraOn catalog). This scarcity for some species can be due to subjective issues like the visual attractiveness of the plant, e.g., having a showy flower, or it can be a real effect, reflecting its rare status in the wild. This effect makes the raw data distribution long-tailed (Figure 3(a), shown in logarithmic scale).
The Floralens dataset was derived by sampling the raw data as follows. First, we filtered out species with less than 50 images. Then, for each of the remaining species, we sampled up to 200 images from the datasets, prioritizing data sources in the following order: (1) FloraOn; (2) Pl@ntNet; (3) Observation.org, and; (4) iNaturalist. That is, for the 50-200 image target per species, we use up as many images as possible from FloraOn first, then from Pl@ntNet, and so on.
(a) Raw data.
(b) Floralens dataset.
Figure 3: Dataset histograms ($x$-axis: number of images; $y$-axis: number of species).
This source-based prioritization intends to define a dataset where images are less prone to identification errors. It takes into account the curation processes associated with each data source. FloraOn is curated by botanists and features high-quality images. These often feature subtle details that help secure the identification of a species. Pl@ntNet data goes through a curation process that involves machine learning, contributors’ reputation scores, and geo-based species verification [46]. Observation.org data can result from automatic validation through image recognition coupled with a check for other approved observations in the geographical vicinity, or through an expert volunteer when automated validation fails [45, 48]. Finally, iNaturalist identifications result from a crowd-sourcing effort whereby a “research-grade” identification for a photo can be obtained from the consensus of just two citizen scientists [44, 49].
We used photos without distinction of plant features as, in general, this metadata was not available from the curated datasets. Similarly, we do not attempt to gauge “photo quality”. The citizen-science platforms provide guidelines for submitting acceptable quality photos, and their curation processes mostly filter inadequate items (e.g., out of focus, under/over-exposure, presence of significant digital noise, insufficient resolution, subject not sufficiently separated from surrounding vegetation).
This procedure resulted in the Floralens dataset, a collection of approximately 300K labeled images covering 1,678 species. As illustrated in Figure 3(b), there are 200 images or very close to it for most of the species. The image count is 200 for 67% (1,128) of the species, 150 or higher for 79% (1,323), and 100 or higher for 86% (1,449). The data source prioritization scheme lead to a more significant fraction of FloraOn images in comparison to the raw data (the fraction grows from 0.5 to 5.1%) and, also, to a relatively even distribution of images from iNaturalist, Observation.org, and Pl@ntnet (the corresponding fractions are 30.7, 29.3, and 34.9%).
4 Model Derivation
The process of deriving an image classification model using GAMLV is illustrated in Figure 4. Overall, it comprises three stages: (1) preparing the data set for training; (2) training the model, and (3) deploying the model onto a cloud server or (using a suitable format) onto edge devices. GAMLV essentially requires the user to focus on the dataset preparation (1), given that training (2) and deployment (3) merely require simple high-level options by the user and are otherwise automated [50, 51]. The interaction with GAMLV can be conducted via a browser with a simple user interface, as we illustrate partially in this section (cf. Figure 5), or programmatically using Google Cloud APIs (e.g., in Python).
Figure 4: Model derivation using GAMLV.
4.1 Preliminaries
The first step requires the user to load the dataset images onto a storage bucket, in this case, provided by the Google Cloud Storage service (GCS), along with a simple CSV file. The latter lists the GCS image URIs and associates each URI to a ground truth label (the name of the species in the image) and to either the train, validation, or test subset.
We fed GAMLV with train, validation, and test splits over the Floralens dataset, corresponding to fractions of 80%, 10%, and 10%. As usual, the train split is used to adjust CNN parameters during the training process in an iterative feedback loop, the validation split is used to measure the progress and convergence of that training process, and the test split is used merely for evaluating the model after training. The splits, with the image counts and data source provenance detailed in Table 2, resulted from a random selection of images for each species. Since the selection process is random and given the volume of images at stake, the overall fraction of images of each data source in each split closely matches that of the overall dataset.
Table 2: Train, validation, and test splits over the Floralens dataset.
Data source Train   Valid.   Test   Total
FloraOn 12,175 1,466 1,550 15,191 ( 5%)
iNaturalist 72,174 8,924 9,029 90,127 (31%)
Observation.org 68,614 8,603 8,529 85,746 (29%)
Pl@ntNet 81,918 10,367 10,252 102,537 (35%)
All 234,881 (80%) 29,360 (10%) 29,360 (10%) 293,601 (100%)
In [14], we consider other strategies for defining these splits. In particular, we explored approaches that prefer specific data sources for the validation/test splits. We found that a random split, besides preserving a roughly similar fraction of images per data source in each split, results in models with better performance (contrast the results in Section 5 with those in [14]). In any case, prioritizing particular data sources (e.g. FloraOn or Pl@ntnet) for validation/test splits over others had little impact on model performance.
(a) Training parameters.
(b) Deployment options.
Figure 5: GAMLV interface for model training and deployment.
4.2 Training
Once the dataset is imported onto AutoML, training may proceed, requiring only the user to make high-level choices for the type of model to be generated and the maximum training time, as illustrated in Figure 5(a). Since we wish to use the model as part of web or mobile applications (cf. Section 6) rather than deploying it on a Google Cloud server, we select the option “Edge” model. We also toggle the option for a model that favors accuracy over latency among the three available choices. The maximum training time is specified as a “node hours” budget, where nodes are virtual machines used during training.
GAMLV required 4 node hours to complete the training of the CNN with the Floralens dataset. The service operates as a “black box”. Thus, it is not possible to discern what goes on during training. For instance, no exact details or configuration options are provided for the training infrastructure (e.g., in terms of virtual machines, GPUs, or TPUs) and it is not possible to track details regarding the training process (e.g., how the model converges over time).
4.3 Deployment
Once training is completed, a model can be deployed in one of several formats amenable for integration with a local application (Figure 5(b)). The formats include the standard SavedModel format used by TensorFlow, but also others like TF Lite [52], a lightweight TensorFlow format for use in resource-constrained hosts (e.g., mobile and embedded devices), or TFJS [53], for use in web browsers or Javascript programs. We use the TF Lite and TFJS variants in the software artifacts described in Section 6.
(a) Initial (including input layer).
(b) Intermediate.
(c) Final (including output layer).
Figure 6: Layers of the CNN model (fragment).
4.4 Derived Model
The model obtained by GAMLV is a 65-layer deep CNN, with the structure partially illustrated in Figure 6 for the TF Lite version, in terms of the input/initial layers (a), intermediate layers (b), and final/output layers (c). Each box in the image represents a type of neuron in the CNN, working as a function that takes a multi-dimensional (e.g., 2D, 3D, …) vector, known as a tensor, and produces an output vector (output tensor). This function is parameterized by internal weights repeatedly adjusted using techniques such as back-propagation as part of an iterative training process. CNNs use a particular family of functions called convolutions, which are especially suited for detecting image features (e.g., edges). In the simplest type of convolution, operating over 2D matrices, each position in the output vector called a feature map, is the dot product of a sliding window over the input matrix with a filter defined by the internal weights of the neuron. For details, see for instance Chapter 9 of [54].
The TF Lite version differs from the standard TensorFlow model only in terms of post-training optimizations like quantization (conversion of floating point weights to an integer scale) that enable the model to be interpreted faster with little degradation in accuracy [52]. As shown in Figure 6, the input layer takes a $224\times 224\times 3$ tensor, corresponding to a $224\times 224$ (typically resized to conform with the dimensions of the CNN input) image with 3 RGB channels, with 8-bit values per color channel. The intermediate layers use several types of convolutions in a repeating pattern. The final layers include the derivation of a 1,280-feature map (a vector summarising all the captured image features [54]). This feature map is then passed to a fully connected soft-max activation function that produces the final classification vector with the label probabilities, 1,678 of them in line with the number of species covered. In this context, fully connected means that every value of the feature map, combined with the internal weights, is taken into account to calculate the value of each output position. The soft-max activation function is used to produce the neural network’s final output which, in this case, is a probability distribution [54]. As expected for a probability distribution, the sum of all the probabilities in the output vector equals 1. Note that we always get such a vector, even for images outside the domain in question (in our case the native Portuguese flora). If the model is certain of a particular species, we get a high probability value for that species and low values for all the others. It can happen of course that the model is less certain, for instance when the probabilities assigned to the most high-ranked species are close (e.g., two or more very similar species). Finally, when the model fails to make any meaningful identification we get low probability values for all species.
The CNN architectures at stake are picked from the MnasNet family [55], developed with mobile and embedded devices in mind. The high-level choice between models offered by GAMLV (back in Figure 5(a)) corresponds to three different MnasNet instantiations that do not differ in structure, just in the density of connections between layers and number of internal weights.
5 Results
We evaluated the Floralens model using the test split described in the previous section, hereafter designated by FLTS (Floralens test split), using standard metrics. We then complemented these baseline results with those obtained using two additional test sets described further in this section [56, 43]. Next, we briefly evaluated the model’s capability of identifying the genus (as opposed to the species), the rationale being that sometimes even if the species cannot be identified with good confidence, knowledge of the genus will be useful to the user. Finally, we compared the Floralens results for all test sets with Pl@ntnet models accessible through the Pl@ntNet API [57].
5.1 Method
The performance of the Floralens model was evaluated over several test datasets using the following metrics. Precision is the ratio of true positives (TP) relative to the total number of positives (TP $+$ FP). A positive (identification) occurs when the classification score (a probability) returned by the model equals or exceeds a confidence level set as the threshold for analysis. Recall is the ratio of true positives relative to the total number of true examples (TP $+$ FN). The confidence level is relevant for practical uses of a model. Using a low confidence level means that false positives are tolerable for the intended use, and we will get lower precision. On the other hand, a too high value for the confidence level will result in more false negatives hence lower recall.
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}\qquad\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$
Top-1 is the fraction of test images that the model correctly classified by the label with rank 1 (the highest-scoring label). Top-5 is similar to Top-1 but accounts for test images with a rank lower or equal to 5 (the 5 highest-scoring labels). We also use a variant of the Mean Reciprocal Rank (MRR) for test images of rank less or equal to 5. These are defined as follows:
[…]
5.2 Baseline Results
In this and the following subsections, larger values indicate better model performance.
(a) Precision-Recall curve.
(b) Values per confidence level.
Figure 7: FLTS results: precision and recall results.
Figure 7 shows the results for precision and recall for the Floralens model applied to the test set given in Table 2, more precisely the macro-average of precision and recall values for all species. The area-under-curve (AUC) for the precision-recall correlation (in 7(a)), also known as the average precision, is 0.72 (the maximum value would be 1.0). Putting the confidence levels in perspective (in 7(b)) we can visualize that precision and recall are approximately equal to 0.7 for a confidence level of 0.2. For a confidence level of 0.5 precision equals 0.85 and recall equals 0.53. Overall, the results indicate a reasonable predictive power for the Floralens model.
Table 3: FLTS results: Top-1, Top-5 and MRR.
Data source Top-1 Top-5 MRR
FloraOn 0.70 0.88 0.77
iNaturalist 0.70 0.87 0.77
Observation.org 0.64 0.83 0.72
Pl@ntNet 0.66 0.87 0.74
Overall 0.67 0.86 0.75
Table 3 lists the Top-1, Top-5, and MRR results for the FLTS, per data source used in the construction of the data set (Floralens, iNaturalist, Observation.org, and Pl@ntNet) and also in overall terms (last line in the table). The results indicate relatively homogeneous predictive power across all data sources, as the maximum difference in values between them does not exceed 0.06: a Top-1 value of 0.64 for Observation.org vs. corresponding values of 0.7 for FloraOn and iNaturalist. The overall measures again indicate reasonably good predictive power: 0.67 for Top-1 (i.e., roughly two-thirds of the FLTS images are correctly classified with rank 1), 0.85 for Top-5, and 0.75 for MRR.
5.3 PlantCLEF and Wikipedia Test Sets
We considered two additional test sets: a random sample of 10,000 labeled images from the PlantCLEF’22-23 [56, 43] competition, and a sample of close to 1,500 images from Wikipedia. The PlantCLEF data we used was only a small sample of the entire “trusted” training set of PlantCLEF [58] that comprises approximately 2.9 million images covering 80,000 plant species. The repository is trusted as the image labels were obtained from academic sources or collaborative platforms like Pl@ntNet or iNaturalist. Our subset was built by first filtering out species not covered by the Floralens model, obtaining data for 1,593 (out of 1,678) species, and then randomly sampling 10,000 images.
As for the Wikipedia test set, the images were identified through the Wikimedia REST API search functionality [59]. For each species in the Floralens domain, we used the species name as the keyword for a REST API search. Among other information, the search result typically yields a reference to an image stored at Wikipedia which we then considered for addition to the test set. After obtaining the images, we filtered out illustrations, herbarium specimens, and duplicates (associated with more than one species). Duplicates typically arise because the search may yield an image of a different species in the same genus, e.g., if the target species’ name does not have a Wikipedia page. Through this process, we obtained a dataset of 1,351 images for an equal number of species (one image per species). Compared to the PlantCLEF test set, the identifications associated with these images are less reliable as they result from an uncontrolled crowd-sourced effort with no specific directives for validation.
Table 4 shows the results of the Floralens model for these test sets considered in terms of the Top-1, Top-5, and MRR metrics. We also recall the overall FTLS results (from Table 3) for easy comparison. We can observe that the PlantCLEF and Wikipedia test set results are marginally lower than those obtained for the FLTS, (by 0.02/0.03 in all metrics). The results show that the Floralens model performs well with datasets other than our base test suite.
Table 4: Top-1, Top-5, and MRR of the Floralens model for all test sets.
Dataset #I #S Top-1 Top-5 MRR
FLTS 29,360 1,678 0.67 0.86 0.75
PlantCLEF 10,000 1,593 0.65 0.84 0.73
Wikipedia 1,351 1,351 0.65 0.84 0.72
5.4 Genus Identification Results
Table 5 shows the results of genus (as opposed to species) identification. The classification score for a genus is obtained by summing the scores for all the species in that genus output by the model. In [14] we created a model specific for genus identification but found that the results were essentially the same as obtained using this aggregation method. The greatest enhancement is observed for the Wikipedia test set, especially, for the Top-1 result ($\Delta=+0.14$). This probably originates from our previous observation that even when the image on the page for a species is wrongly labeled, Wikipedia does manage to provide an image of a plant of the same genus. The improvements observed for FLTS and PlantCLEF for all metrics are the same.
Table 5: Top-1, Top-5, and MRR of Floralens for genus prediction ($\Delta$: variation relative to species results).
Dataset Top-1 $\Delta$ Top-5 $\Delta$ MRR $\Delta$
FLTS 0.76 +0.09 0.91 +0.05 0.82 +0.07
PlantCLEF 0.74 +0.09 0.89 +0.05 0.80 +0.07
Wikipedia 0.79 +0.14 0.91 +0.07 0.83 +0.08
5.5 Comparative Pl@ntNet API Results
We now provide results comparing the Floralens model with models accessible via the Pl@ntNet API [57]. The Pl@ntNet API is a web service providing access to the same visual identification models used by state-of-the-art Pl@ntNet apps [6]. The API lets us obtain a set of ranked species for a given image for two models for worldwide flora: a so-called “legacy” model from 2022 (henceforth ${\rm PN}^{22}$) generated using CNN, and; a recent model announced in July 2023 [19], generated using Vision Transformers, henceforth ${\rm PN}^{23}$. Through the API, it is also possible to filter results from the ${\rm PN}^{23}$ model so that only species occurring in a specific biogeographic region are included. One such region is Southwestern Europe including Portugal, allowing the most head-to-head comparison between Floralens and Pl@ntNet possible. These results are identified by ${\rm PN}^{23\text{F}}$.
Table 6: Pl@ntNet API: comparative MRR values ($\Delta$: variation relative to the Floralens model).
Dataset ${\rm PN}^{22}$ $\Delta$ ${\rm PN}^{23}$ $\Delta$ ${\rm PN}^{23\text{F}}$ $\Delta$
FLTS 0.68 --0.07 0.80 +0.05 0.80 +0.05
PlantCLEF 0.72 --0.01 0.79 +0.06 0.79 +0.06
Wikipedia 0.73 +0.01 0.78 +0.06 0.79 +0.07
Table 7: Floralens vs Pl@ntNet API: MRR per data source in the FLTS ($\Delta$: variation relative to the Floralens model).
Source ${\rm PN}^{22}$ $\Delta$ ${\rm PN}^{23}$ $\Delta$ ${\rm PN}^{23\text{F}}$ $\Delta$
FloraOn 0.58 --0.17 0.79 +0.04 0.79 +0.04
iNaturalist 0.67 --0.10 0.81 +0.04 0.81 +0.04
Observation.org 0.59 --0.13 0.76 +0.04 0.76 +0.04
Pl@ntNet 0.77 +0.03 0.84 +0.10 0.84 +0.10
[…]
FLTS 0.68 --0.07 0.80 +0.05 0.80 +0.05
Table 6 shows the variation ($\Delta$) of the MRR values obtained for ${\rm PN}^{22}$, ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$ for all the test sets relative to the corresponding values obtained for Floralens. ${\rm PN}^{22}$ performs worse than Floralens for the FLTS, a variation of $-0.07$. The MRR values of ${\rm PN}^{22}$ are otherwise similar for PlantCLEF ($-0.01$) and Wikipedia ($+0.01$). The discrepancy observed for FLTS merits further analysis and is discussed below. Focusing now on ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$, the MRR values across all test sets range from 0.78 to just 0.80, and perform better than Floralens by a factor of 0.05 to 0.07. The Southwestern Europe species filter associated with ${\rm PN}^{23\text{F}}$ has little impact on the results.
[…]
Overall, the Floralens results are on par and in some cases better than ${\rm PN}^{22}$, and marginally worse than ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$.
6 Software Artifacts
In this section, we briefly describe the Biolens software platform, in which the Floralens has been made publicly available, as well as other artifacts provided to the community.
6.1 Biolens Website
The Floralens model has been integrated into the Biolens project website [9]. As illustrated in the screenshots of Figure 8, the functionality is quite simple: (a) users submit photos of interest, and (b) obtain suggestions for identifications, with a textual indication of the model’s confidence. For instance, a “HIGH” confidence label as shown in (b) means that the model has a score higher than 70% for the species at stake; in other cases, we can have “MEDIUM” or “LOW” confidence labels corresponding to a score of 40-70% or lower than 40%, respectively. Up to 5 species are listed by the app, as long as the model outputs a minimum score of 15% for each identification. The Biolens website is hosted by a small virtual machine that requires just 2 CPU cores and 8 GB of RAM. With this configuration, invoking the model takes an average of 900 milliseconds per image, as calculated directly from approximately 28,000 requests to the server between April 2021 and September 2024. As for the model itself, it takes only 8 MB of disk storage. The server configuration is lightweight as we use the TF Lite variant of the Floralens model running on the server side (similarly to the other models for other taxa also hosted on the site). The user’s browser merely displays the server’s results, it does not host the model.
(a) Image uploading by the user.
(b) Suggested automatic identification.
Figure 8: Biolens – web application screenshots.
6.2 Biolens App
We also recently developed a prototype version of a mobile application that can run on Android and iOS devices. The Android version is available for download at the Biolens website. A few screenshots of the application are shown in Figure 9. The functionality is similar to that of the Biolens website, but customized for a mobile application context: users can take photos of specimens on the fly and obtain instant identification suggestions without an Internet connection. All Biolens models are bundled within the app and, thus, are evaluated in loco on the mobile device. The identification information, the date, the current geographical location, and optional user annotations, are recorded in association with each photo. According to the reference latency values provided by Google AutoML for devices like Google Pixel 2, Samsung Galaxy S7, or iPhone X, invoking the model takes less than 100 milliseconds or lower in modern mobile devices. This is consistent with our experience and much faster than in the case of the Biolens server discussed above. The server’s configuration is quite modest compared to present-day mobile devices.
(a) Model selection.
(b) Image list.
(c) Details for an image.
Figure 9: Biolens – mobile application screenshots.
6.3 Dataset and Results
The full Floralens dataset is publicly available on Zenodo [16]. The dataset contains the mapping between the image labels (ground truth), the image URLs from which they were retrieved, URLs for a site we maintain where all images are also stored, and GBIF identifiers when applicable (all images except those obtained from FloraOn). Ground truth and URLs are also available for the PlantCLEF and Wikipedia datasets used in the evaluation of Section 5, as well as for all datasets the Top-5 results and corresponding confidence levels for the Floralens model and the three Pl@ntNet model variations.
7 Conclusions
In this paper, we describe the construction of a dataset for the Portuguese flora and the derivation of a deep-learning model for the automatic identification of the species therein. The universe of species was taken from the FloraOn dataset, provided by the Sociedade Portuguesa de Botânica and compiled exclusively by specialists. The dataset was constructed based on high-quality data from several research-grade datasets available via GBIF. Besides FloraOn these include: iNaturalist, Pl@ntNet, and Observation.org. We made the dataset available to the community on Zenodo [16]. The Floralens model was derived from this dataset using GAMLV which provides users with tools to derive models from datasets using off-the-shelf convolutional deep neural networks. The model is available online at the BioLens Project website and can also be used as part of the Biolens mobile application (e.g., offline, in the field).
The Floralens model has good predictive power, with an AUC metric value of 0.72 and, for a reference confidence level of 0.5, values for precision and recall of 0.85 and 0.53, respectively. It also features a relatively homogeneous predictive power across all data sources used in the dataset, with a maximum variation of 0.06, and values for Top-1 and Top-5 of 0.67 and 0.86, respectively. Compared with the state-of-the-art platform Pl@ntnet, Floralens performed on par with the “legacy 2022” model and only marginally worse when compared with the most recent model.
As for future work, we aim to improve the species coverage and the model’s accuracy. One way to do that is to include data from other datasets such as those of Encyclopedia of Life [60] and FloraIncognita [5]. With hindsight, also including images from Morocco and Algeria would probably have helped enhance the dataset with images from species in common with the Mediterranean flora.
We also want to address limitations that arise from the inconsistent use of taxonomic names and synonyms. Our list of Portuguese native species taken from FloraOn is as complete and up-to-date as possible. Recent taxonomic revisions changed the binomial names for some species. This is not necessarily reflected in the public datasets. For example, the species featured in the FloraOn listing as Atractylis gummifera is now known as Chamaeleon gummifer. Although not ubiquitous, it is widespread in the Mediterranean region [61]. Nevertheless, it is not included in the Floralens dataset as there weren’t enough ($\geq$ 50) photos available in GBIF. However, a recent query for Chamaeleon gummifer yields more than enough photos to include the species in the Floralens dataset in a future update.
More work is also required on the Biolens mobile app to improve its usability and optimize resource usage. Integrating with existing Citizen Science platforms is one possibility, allowing the user to upload Biolens records.
Finally, we plan to continue preliminary work on the use of image similarity models [14]. These may provide an alternative way to clinch an identification when GAMLV-based models yield low-confidence results. Hybrid classification models that combine both approaches are an interesting possibility.
Acknowledgements.
The authors would like to thank Hugo Gresse, Pierre Bonnet, and Mathias Chouet from Project Pl@ntNet for kindly providing us with extended access to the Pl@ntNet API for our analysis. This work was partially funded by projects SafeCities and Augmanity (POCI-01-0247-FEDER-041435 and -046103, through COMPETE 2020 and Portugal 2020), and by project UIDB/50014/2020 (Fundação para a Ciência e Tecnologia). It would not have been possible without the support from the Google Cloud Research Credits program. Finally, we thank the anonymous referees whose comments allowed us to improve this paper.